# Bourbaki degree of pairs of projective surfaces

### v2

@@ 3.2 The Bourbaki degree of a plane curve from the Bourbaki degree of a pair of surfaces
Let us denote by $\nabla\overline{g}=(\partial_{0}g,\partial_{1}g,\partial_{2}g)$ the vector, so we denote by
@@
the kernel of the multiplication by this vector. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
@@
and by the snake lemma, we obtain isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$.
@@
$\begin{cases}c_{2}({\mathcal{T}}_{\sigma}(e))&=\operatorname{Bour}(\sigma)=e(e-d_{g})+d_{g}^{2}-m(\sigma)\\ c_{2}(\ker\nabla g(e))&=\operatorname{Bour}(X)=e(e-d_{g})+d_{g}^{2}-\tau(X),\end{cases}$
@@
From the construction, we have that the singular scheme of ${\mathcal{T}}_{\sigma}$ (described by $Z_{g}$) intersects $V(x_{3})$ transversely, so we obtain the formula between Chern classes $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:\mathbb{P}^{2}\simeq V(x_{3})\hookrightarrow\mathbb{P}^{3}$ is the inclusion. Therefore,
@@
and in particular $\operatorname{Bour}(\sigma)=\operatorname{Bour}(X)$.
Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number, and thus it should be the correct generalization for the case of pairs of surfaces.
3.3 Locally free resolutions and the Bourbaki scheme
In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [17, Theorem 2.1, (c)]. This follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$.
@@ 3.3 Locally free resolutions and the Bourbaki scheme
The 0-th Fitting ideal coincides with the annihilator ideal of $\mathcal{Q}_{\sigma}$, and the $1$-st Fitting ideal has dimension zero.
The minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0,$
so that $e=1$, $m(\sigma)=4$ and $\operatorname{Bour}(\sigma)=1$, with $c_{3}({\mathcal{T}}_{\sigma})=3$.
@@
$\nabla\sigma=\begin{pmatrix}x_{2}x_{3}&-x_{2}x_{3}&x_{0}x_{3}-x_{1}x_{3}&x_{0}x_{2}-x_{1}x_{2}\\ 3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}&2x_{0}x_{1}&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$.
In this case, the 0-th Fitting ideal coincides with the annihilator ideal of $\operatorname{coker}(\nabla\sigma)$, and the saturation of the first Fitting ideal is zero. The primary decomposition of $\Xi_{\sigma}$ is described in the following table:
dimension degree radical ideal
1 1 $(x_{2},x_{3})$
1 1 $(x_{3},x_{0}-x_{1})$
1 1 $(x_{2},x_{0}-x_{1})$
1 2 $(x_{0},x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$
0 4 $(x_{2}-x_{3},x_{0}x_{1}-x_{1}^{2}+x_{3}^{2},x_{0}^{2}+x_{1}^{2})$
0 4 $(x_{2}-x_{3},x_{0}x_{1}-x_{1}^{2}+x_{3}^{2},x_{0}^{2}+x_{1}^{2})$
@@
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for $\mathcal{I}_{B}$ will be of the form:
@@
    if $\sigma$ is nearly-free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
@@
since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for the syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. ∎
@@
The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [5], related to rational cuspidal curves. In [17, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [5] (see [17, Definition 2.17]). Here, we are inspired by their definition, since this is equivalent to the notion using the minimal free resolution.
The notion of 33-syzygy divisors is also present in a number of works in the field, for example [1], [6] and [8].
@@
  • For $k>2$, $(f,g)$ is nearly-free with $e=1$.
@@
and in particular $\sigma$ is a nearly-free sequence.
@@
and focus on the second one. The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [14, Proposition 1.1]. Furthermore, since $\operatorname{{rk}}(F_{0}^{\prime})=3$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence
@@
The following proposition follows from the formula in Remark 7 for $c_{3}({\mathcal{T}}_{\sigma})$ in terms of $e,\deg(B)$ and $p_{a}(B)$ for a Bourbaki scheme $B=B_{\nu}$.
Proposition 21.
If $\sigma=(f,g)$ is a nearly-free sequence such that ${\mathcal{T}}_{\sigma}$ is locally free, then $d=d_{f}+d_{g}$ must be even and ${\mathcal{T}}_{\sigma}$ is $\mu$-stable.
Proof.
Since
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e),$
assuming ${\mathcal{T}}_{\sigma}$ is locally free, we obtain $c_{3}({\mathcal{T}}_{\sigma})=0$. On the other hand, since $\sigma$ is nearly-free, $\deg(B)=1$ and $p_{a}(B)=0$, hence
$e=\frac{d+2}{2},$
which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$. ∎

### v3

@@ 2.2 Plane curves as pairs of surfaces
Let us denote by $\nabla\overline{g}=(\partial_{0}g,\partial_{1}g,\partial_{2}g)$ and
@@
the kernel of the multiplication. Using the block-form of the matrix $\nabla\sigma$, may form the following diagram with exact columns:
@@
and by the snake lemma, we obtain isomorphisms ${\mathcal{T}}_{\sigma}\simeq{\mathcal{T}}_{\overline{g}}$ and $\mathcal{Q}_{\sigma}\simeq\mathcal{O}_{Z_{g}}(d)$. Denoting $i:H\hookrightarrow\mathbb{P}^{2}$ the inclusion, it is clear that $i^{*}({\mathcal{T}}_{\sigma})\simeq{\mathcal{T}}_{g}$ as a logarithmic sheaf over $\mathbb{P}^{2}$, since $H$ and $Z_{g}$ intersect transversely.
@@
$\begin{cases}c_{2}({\mathcal{T}}_{\sigma}(e))&=\operatorname{Bour}(\sigma)=e(e-d_{g})+d_{g}^{2}-m(\sigma)\\ c_{2}({\mathcal{T}}_{g}(e))&=\operatorname{Bour}(X)=e(e-d_{g})+d_{g}^{2}-\tau(X),\end{cases}$
@@
From the transversality, we obtain $i^{*}(c({\mathcal{T}}_{\sigma}))=c(i^{*}({\mathcal{T}}_{\sigma}))$ where $i:H\hookrightarrow\mathbb{P}^{3}$ and thus
@@
and in particular $\operatorname{Bour}(\sigma)=\operatorname{Bour}(X)$. Geometrically, this means that $m(\sigma)$ counts the singular lines of the ruled surface $S=V(g)\subset\mathbb{P}^{3}$ with the same multiplicity as the Tjurina number does.
A more algebraic way to see this is through the associativity formula. Note that $\mathcal{Q}_{\sigma}\otimes\mathcal{O}_{H}\simeq\mathcal{O}_{\operatorname{Sing}(X)}(d)$, with the effect of adding a variable to the graded ring of the plane $\mathbb{P}^{2}\hookrightarrow\mathbb{P}^{3}$, and the associated map of modules is given by the quotient map
$0\to(x_{3})\to\kappa[x_{0},x_{1},x_{2},x_{3}]\to\kappa[x_{0},x_{1},x_{2}]\to 0.$
Then, for any associated prime $P$ of the Jacobian ideal of $X$ at $\kappa[x_{0},x_{1},x_{2}]$, the extension will satisfy
$\dim R/(R\cdot P)$ $=\dim(\kappa[x_{0},x_{1},x_{2}]/P)+1$
$\deg R/(R\cdot P)$ $=\deg(\kappa[x_{0},x_{1},x_{2}]/P).$
since the generators do not include $x_{3}$. Moreover, the length of the modules over $P$ will be preserved, as the element $x_{3}\not\in P$ will be invertible. Together with the associativity formula, we obtain:
$\tau(X)$ $=\sum_{P\in\operatorname{Ann}(\mathcal{O}_{\operatorname{Sing}X}(d)),\dim(\kappa[x_{0},x_{1},x_{2}]/P)=1}\operatorname{{length}}(\mathcal{O}_{\operatorname{Sing}X}(d))_{P}\cdot\deg(\kappa[x_{0},x_{1},x_{2}]/P)$
$=\sum_{P\in\operatorname{Ann}(\mathcal{Q}_{\sigma}),\dim(\kappa[x_{0},x_{1},x_{2},x_{3}]/P)=2}\operatorname{{length}}(\mathcal{Q}_{\sigma})_{P}\cdot\deg(R/P)=m(\sigma).$
The following example of the nature above is a sequence $\sigma=(x_{3},g)$ of degree $d$ such that $e=d$ but $m(\sigma)\neq 0$, showing that a converse for 8, $(a)$ is not possible without assuming the Buchsbaum–Rim resolution.
Example 11.
Let $f=x_{3}$ and $g=x_{0}x_{1}x_{2}^{2}+x_{0}^{4}+x_{1}^{4}$. The quartic plane curve $X=V(g)\subset\mathbb{P}^{3}_{[x_{0}:x_{1}:x_{2}]}$ is nodal and the minimum degree for a syzygy of $\nabla g$ is $e=d$.
Hence, $m(\sigma)=\tau(X)=1\neq 0$, accounting for the nodal singularity, but $e=d=3$.
2.3 Free resolutions and the Bourbaki scheme
In this section, we relate resolutions for $B$ and for ${\mathcal{T}}_{\sigma}$, and we use this relationship to characterize sequences, in the spirit of [undefr, Theorem 2.1, (c)]. The next result follows analogously since $H^{1}(\mathcal{O}_{\mathbb{P}^{3}}(*))=0$, as it is the case for $\mathbb{P}^{2}$.
@@ 2.3 Free resolutions and the Bourbaki scheme
The scheme $(\Xi_{\sigma})_{1}$ has three primary components in dimension one, given by the prime ideals $\mathfrak{p}_{1}=(x_{0},x_{3})$, $\mathfrak{p}_{2}=(x_{0},x_{2})$ and $\mathfrak{p}_{3}=(x_{0},x_{1})$. We also note that there is an easy syzygy of degree one, namely $\nu=(-x_{3},0,0,x_{0})^{T}$, hence $e=1$.
To show $m(\sigma)=4$, we study the length of the cokernel module $Q$ associated to the sheaf $\mathcal{Q}_{\sigma}$ over each prime $\mathfrak{p}_{i}$, $i=1,2,3$, and use the associativity formula.
Over $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that $x_{1}x_{2}\in R_{P}^{\times}$ is a unit, and $3x_{0}^{2}\in\mathfrak{p}_{P}$ is in the maximal ideal, hence $v=3x_{0}^{2}+x_{1}x_{2}$ is also invertible, denote $u=v^{-1}$. From this, we may rewrite the matrix as
$(\nabla\sigma)_{\mathfrak{p}_{1}}=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 1&ux_{0}x_{2}&ux_{0}x_{1}&3ux_{3}^{2}\\ \end{pmatrix}\sim\begin{pmatrix}0&-2x_{0}^{2}x_{2}u&-2ux_{0}^{2}x_{1}&2x_{3}-6ux_{0}x_{3}^{2}\\ 1&x_{0}x_{2}u&x_{0}x_{1}u&3x_{3}^{2}u\\ \end{pmatrix}$,
first by multiplying the second row by $u$, and then adding the second row scaled by $2x_{0}$ to the first row. From the final form of the matrix above, it is easy to see that it sends the fourth basis vector to $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, and thus we may compute the cokernel as the remaining entries of the first row:
$\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{1}}\simeq\frac{R_{\mathfrak{p}_{1}}}{(-2x_{0}^{2}x_{2}u,-2ux_{0}^{2}x_{1},2x_{3}-6ux_{0}x_{3}^{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{0}^{2},x_{3})},$
a $R_{\mathfrak{p}_{1}}$-module of length two.
Over the prime $\mathfrak{p}_{2}=(x_{0},x_{2})$, the element $x_{3}\in R_{\mathfrak{p}_{2}}$ is an unit. We denote by $u=x_{3}^{-1}\in R_{\mathfrak{p}_{2}}$ its inverse, and we rewrite the matrix using elementary operations
$(\nabla\sigma)_{\mathfrak{p}_{2}}=\begin{pmatrix}2x_{0}&0&0&2x_{3}\\ 3x_{0}^{2}+x_{1}x_{2}-3x_{0}x_{3}&x_{0}x_{2}&x_{0}x_{1}&0\\ \end{pmatrix}\sim\begin{pmatrix}ux_{0}&0&0&1\\ 3x_{0}^{2}+x_{1}x_{2}-3x_{0}x_{3}&x_{0}x_{2}&x_{0}x_{1}&0\\ \end{pmatrix}$,
first by adding the first row scaled by $(-\frac{3}{2}x_{3})$ to the second row, and then multiplying the first row by $u/2$. Now, it is clear that the matrix sends the fourth basis vector to $(1,0)\in R_{P}^{2}$, and thus
$\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0}x_{1},x_{0}x_{2},3x_{0}^{2}+x_{1}x_{2}-x_{0}x_{3})}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{0}x_{2},3x_{0}^{2}+x_{1}x_{2}-x_{0}x_{3})}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{0}x_{2},x_{1}x_{2})}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{0},x_{2})}\simeq\kappa$
since $x_{1}\in R_{\mathfrak{p}_{2}}^{\times}$. Thus, it follows $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=1$. For the ideal $\mathfrak{p}_{3}$, we note that the approach is analogous as the previous one, using the same elementary operations, and obtaining the same isomorphism at the end $\operatorname{coker}(\nabla\sigma)_{\mathfrak{p}_{3}}\simeq\kappa$, so that from the associativity formula:
$m(\sigma)=\sum_{i=1}^{3}\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{i}})\cdot\deg(R/\mathfrak{p}_{i})=2+1+1=4.$
Using the Bourbaki formula with $m_{0}=d_{f}^{2}+d_{g}^{2}+d_{f}d_{g}=7$, $d=3$ and the obtained discrete invariants $m(\sigma)=4$ and $e=1$, we conclude that $\operatorname{Bour}(\sigma)=1$.
From the minimal free resolution of the ideal of a line in $\mathbb{P}^{3}$, using Lemma 12, the minimal free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-4)\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 2}\oplus\mathcal{O}_{\mathbb{P}^{3}}(-1)\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
From the resolution, one obtains $c_{3}({\mathcal{T}}_{\sigma})=3$.
@@
$\nabla\sigma=\begin{pmatrix}x_{2}x_{3}&-x_{2}x_{3}&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\\ 3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}&2x_{0}x_{1}&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$.
The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines and a plane quadric, given by the following prime ideals of $R$:
$\mathfrak{p}_{1}$ $=(x_{2},x_{3})$
$\mathfrak{p}_{2}$ $=(x_{3},x_{0}-x_{1})$
$\mathfrak{p}_{3}$ $=(x_{2},x_{0}-x_{1})$
$\mathfrak{p}_{4}$ $=(x_{0},x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$
We claim that $m(\sigma)=5$, a quantity which will be obtained using the associativity formula. For $\mathfrak{p}_{1}=(x_{2},x_{3})$, we start by noting that the quantities:
$h$ $=3x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2}$
$v$ $=x_{0}x_{1}$
$t$ $=(x_{0}-x_{1})$
are all invertible inside the ring $R_{\mathfrak{p}_{1}}$, where the first one is the sum of an invertible element with an element inside the maximal ideal $(x_{2},x_{3})_{\mathfrak{p}_{1}}$. Then, we do two elementary operations:
$(\nabla\sigma)_{\mathfrak{p}_{1}}$ $\sim\begin{pmatrix}ux_{2}x_{3}+\frac{1}{2}hx_{2}x_{3}&0&utx_{3}+x_{0}x_{2}^{2}x_{3}&utx_{2}+x_{0}x_{2}x_{3}^{2}\\ h&2v&2x_{0}x_{2}&2x_{0}x_{3}\end{pmatrix}$
$\sim\begin{pmatrix}ux_{2}x_{3}+\frac{1}{2}hx_{2}x_{3}&0&utx_{3}+x_{0}x_{2}^{2}x_{3}&utx_{2}+x_{0}x_{2}x_{3}^{2}\\ \frac{1}{2}uh&1&ux_{0}x_{2}&ux_{0}x_{3}\end{pmatrix}$.
The first is to add $\frac{1}{2}x_{2}x_{3}$ times the second row to the first row, scaled by $u=v^{-1}$. Then, we scale the second row by $\frac{1}{2}u$. Since the matrix sends the second basis vector into $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, it suffices to look at the first row of the remaining matrix, so that
$\operatorname{coker}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{((u+\frac{1}{2}h)x_{2}x_{3},utx_{3}+x_{0}x_{2}(x_{2}x_{3}),utx_{2}+x_{0}x_{3}(x_{2}x_{3}))}$
$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2}x_{3},utx_{3}+x_{0}x_{2}(x_{2}x_{3}),utx_{2}+x_{0}x_{3}(x_{2}x_{3}))}$
$\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2}x_{3},utx_{3},utx_{2})}\simeq\frac{R_{\mathfrak{p}_{1}}}{(x_{2},x_{3})}\simeq\kappa,$
and hence $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
For $\mathfrak{p}_{2}=(x_{3},x_{0}-x_{1})$, note that $v=x_{0}x_{2}\in R_{\mathfrak{p}_{2}}$ is a unit and denote $u=v^{-1}$. We consider some elementary operations:
$(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\sim\begin{pmatrix}x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})h&-x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{1}&0&x_{2}(x_{0}-x_{1})-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{3}\\ h&2x_{0}x_{1}&2v&2x_{0}x_{3}\end{pmatrix}$
$\sim\begin{pmatrix}x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})h&-x_{2}x_{3}-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{1}&0&x_{2}(x_{0}-x_{1})-\frac{1}{2}ux_{3}(x_{0}-x_{1})x_{0}x_{3}\\ \frac{1}{2}uh&ux_{0}x_{1}&1&ux_{0}x_{3}\end{pmatrix}$.
The first transformation is to add the scaled second row (by $\frac{1}{2}x_{3}(x_{0}-x_{1})u$) to the first one, and the second transformation is rescaling the second row by $\frac{1}{2}u$. Then, we see that the matrix sends the third basis vector to $(0,1)\in R_{\mathfrak{p}_{2}}^{2}$, so we may compute the cokernel as the first row of the remaining matrix:
$\operatorname{coker}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3}(x_{2}-\frac{1}{2}u(x_{0}-x_{1})h),-x_{3}(x_{2}+\frac{1}{2}u(x_{0}-x_{1})x_{0}x_{1}),(x_{0}-x_{1})(x_{2}-\frac{1}{2}ux_{0}x_{3}^{2})))}$
$\simeq\frac{R_{\mathfrak{p}_{2}}}{x_{3},-x_{3},(x_{0}-x_{1})}\simeq\frac{R_{\mathfrak{p}_{2}}}{(x_{3},x_{0}-x_{1})}\simeq\kappa,$
since for example the quantity
$x_{2}-\frac{1}{2}u(x_{0}-x_{1})h$
is of the form a unit $x_{2}$ plus something inside the maximal ideal $(x_{3},x_{0}-x_{1})_{\mathfrak{p}_{2}}$, it is also a unit. Thus, we obtain $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}})=1$. An analogous consideration also holds for the case of the prime $\mathfrak{p}_{3}=(x_{2},x_{0}-x_{1})$, with similar elementary operations, now to trivialize the fourth column as $(0,1)$, and then to show that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\kappa$.
For the prime $\mathfrak{p}_{4}=(x_{0},h\doteq x_{0}^{2}+x_{1}^{2}+x_{2}^{2}+x_{3}^{2})$, we note that $v=x_{2}x_{3}$ is invertible in $R_{\mathfrak{p}_{4}}$, denote $u=v^{-1}$ and do the elementary operations
$(\nabla\sigma)_{\mathfrak{p}_{4}}$ $\sim\begin{pmatrix}v&-v&x_{3}(x_{0}-x_{1})&x_{2}(x_{0}-x_{1})\\ h+2x_{0}x_{1}&0&2x_{0}x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})&2x_{0}x_{2}x_{3}(x_{0}-x_{1})\\ \end{pmatrix}$
$\sim\begin{pmatrix}-1&1&-ux_{3}(x_{0}-x_{1})&-ux_{2}(x_{0}-x_{1})\\ h+2x_{0}x_{1}&0&2x_{0}x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})&2x_{0}x_{2}x_{3}(x_{0}-x_{1})\\ \end{pmatrix}$,
first adding the scaled first row $(by2x_{0}x_{1}u)$ to the second row, and then scaling the first row by u-u. The matrix sends the second basis vector to $(1,0)\in R_{\mathfrak{p}_{4}}^{2}$, and thus to compute the cokernel we may consider only the second row of the remaining matrix, so that
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}$ $\simeq\frac{R_{\mathfrak{p}_{4}}}{(h+2x_{0}x_{1},x_{0}(2x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1})),2x_{0}x_{2}x_{3}(x_{0}-x_{1}))}$
$\simeq\frac{R_{\mathfrak{p}_{4}}}{(h+2x_{0}x_{1},x_{0},x_{0})}$
$\simeq\frac{R_{\mathfrak{p}_{4}}}{(x_{0},h)}\simeq\kappa,$
noting that the factors $(2x_{2}+2x_{0}x_{1}x_{3}(x_{0}-x_{1}))$ and $2x_{2}x_{3}(x_{0}-x_{1})$ are units in the local ring. Thus, we obtain that the length at each prime is one, but one of them has multiplicity two, therefore $m(\sigma)=5$ from the associativity formula. The main result of the next section is:
$e=1$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 2$
$e=2$ $\Rightarrow\operatorname{Bour}(\sigma)\leq 5,$
so for $m_{0}=12$ we conclude that these two correspond to $m(\sigma)\geq 7$ and $m(\sigma)\geq 6$, respectively. Since $m(\sigma)=5$, we conclude automatically that $e\geq 3$.
Using Macaulay2, we obtain a minimal free resolution for ${\mathcal{T}}_{\sigma}$ of the 33-syzygy kind, given by:
@@
so $e=3$ and, with $m(\sigma)=5$, we conclude $\operatorname{Bour}(\sigma)=4$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. A resolution for the Bourbaki ideal $\mathcal{I}_{B}$ will be of the form:
@@
    if $\sigma$ is nearly free, then the isolated zeros of $(2\times 2)$-minors of $\nabla\sigma$ are aligned.
@@
since $\operatorname{Bour}(\sigma)=1$ if and only if $B_{\nu}$ is a line for a choice of syzygy of minimum degree $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$. The isolated zeros of the minors will be in the support of the sheaf $\operatorname{supp}({\mathcal{T}}_{\sigma})$, and these will be inside the line $l$ from the resolution obtained. ∎
@@
The notion of nearly free curves for plane curves $V(f)\subset\mathbb{P}^{2}$ is first introduced by [undefd], related to rational cuspidal curves. In [undefr, Proposition 2.18], the authors show that $\operatorname{Bour}(f)=1$ if and only if $V(f)\subset\mathbb{P}^{2}$ is a nearly free curve in the sense of [undefd] (see [undefr, Definition 2.17]). Here, we are inspired by their definition, since this is equivalent to the notion using the minimal free resolution.
The notion of 33-syzygy divisors is also present in a number of previous works, for example [undef], [undefe] and [undefg].
@@
  • For $k>2$, $(f,g)$ is nearly free with $e=1$.
The Jacobian matrix for $\sigma=(f,g)$ is
$\nabla\sigma=\begin{pmatrix}3x_{0}^{2}+x_{1}x_{2}&x_{0}x_{2}&x_{0}x_{1}&3x_{3}^{2}\\ (k+1)x_{0}^{k}&0&0&(k+1)x_{3}^{k}\end{pmatrix}$.
@@
and in particular $\sigma$ is a nearly free sequence.
Proposition 19.
If $\sigma=(f,g)$ is a nearly free normal sequence, then ${\mathcal{T}}_{\sigma}$ is not locally free.
Proof.
Since
$c_{3}({\mathcal{T}}_{\sigma})=2p_{a}(B)-2+\deg(B)(4+d-2e),$
assuming ${\mathcal{T}}_{\sigma}$ is locally free, we obtain $c_{3}({\mathcal{T}}_{\sigma})=0$. On the other hand, since $\sigma$ is nearly free, $\deg(B)=1$ and $p_{a}(B)=0$, hence
$e=\frac{d+2}{2},$
which implies both that $d$ must be even (otherwise $e$ is not an integer) and that $e=d/2+1$, giving that $h^{0}({\mathcal{T}}_{\sigma}(l))=0$ whenever $l\leq d/2=-\mu({\mathcal{T}}_{\sigma})$, hence ${\mathcal{T}}_{\sigma}$ is $\mu$-semistable.
If we denote by $E\doteq{\mathcal{T}}_{\sigma}(d/2)$ the normalized vector bundle, we must have $c_{1}(E)=0$ and
$c_{2}(E)$ $=c_{2}({\mathcal{T}}_{\sigma})-\frac{d^{2}}{2}+\frac{d^{2}}{4}$
$=m_{0}-m(\sigma)-\frac{d^{2}}{4}.$
On the other hand, using the Bourbaki degree formula with $\operatorname{Bour}(\sigma)=1$ and $e=\frac{d+2}{2}$, we obtain
$1=\operatorname{Bour}(\sigma)$ $=\frac{(d+2)^{2}}{4}-\frac{d(d+2)}{2}+m_{0}-m(\sigma)$
$=\frac{d^{2}}{4}-\frac{d^{2}}{2}-d+\frac{4d}{4}+1+m_{0}-m(\sigma),$
which yields
$m_{0}-m(\sigma)-\frac{d^{2}}{4}=0,$
so that $c_{2}(E)=0$. But a $\mu$-semistable reflexive sheaf with $c_{1}(E)=c_{2}(E)=0$ must be $E\simeq\mathcal{O}_{\mathbb{P}^{3}}^{\oplus 2}$ ([undefn, Lemma 9.7]), implying $\sigma$ is free, contradicting the hypothesis. ∎
@@
and focus on the second one. The sheaf $S$ is the kernel of a map between a locally free sheaf $F_{0}^{\prime}$ and a torsion-free sheaf ${\mathcal{T}}_{\sigma}$, thus $S$ is reflexive, from [undefn, Proposition 1.1]. Furthermore, since $\operatorname{{rk}}(F_{0}^{\prime})=3$ and $\operatorname{{rk}}({\mathcal{T}}_{\sigma})=2$, $S$ is a reflexive sheaf of rank one, thus $S\simeq\mathcal{O}_{\mathbb{P}^{3}}(-k)$ for some $k\in\mathbb{Z}$, hence